can i please have help i need to find the x intercepts
y=x^2+5x+6

Answer by praseenakos@yahoo.com(507)

QUESTION:

find the x intercepts
y=x^2+5x+6


To find the x intercept of the given equation, put y = 0 in the given equation.


==> 0 = x^2 + 5x + 6

Which is a quadratic equation which you can solve using either quadratic formula ya splitting the middle term

Here we shall use the second method, that is splitting the middle term


x^2 + 5x + 6 = 0


==> x^2 + 3x + 2x + 6 = 0


==> (x^2 + 3x ) + ( 2x + 6 ) = 0


==> x(x+3) + 2(x+3) = 0


==> (x+ 3) (x + 2) = 0



==> x+ 3 = 0 or x = x+2



==> x = -3 or x = -2



So we got two values corresponding to y = 0

So x intercepts of the the given equation are, x = -3 and x = -2




Hope you understood.


Regards.


Praseena.